OpenStudy (anonymous):

how do i create a geometric sequence if someone is taking 500mg of a drug every 8 hours and has to do so for 10 days. every 8 hours, their kidneys eliminate 65% of this drug. ill be completing a table for the amount of drug in the body just after each 500mg does has been given

OpenStudy (anonymous):

if it helps, the equation that shows the inital dose remaining: \[A = 500e ^{-0.131t}\]

OpenStudy (anonymous):

After first dose:\[A_1=500\] After second those:\[A_2=A_1·(1-0.65)+500=500·(1-0.65)+500\] After third dose:\[A_3=A_2·(1-0.65)+500=500·(1-0.65)^2+500+500·(1-0.65)+500\] In general, after the n-th intake:\[A_n=500·[0.35^{n-1}+0.35^{n-2}+... +0.35^1+0.35^0]\]This can be expressed as: \[A_n=500\sum_{k=0}^{n-1}0.35^k\] and 0.35^k is a geometric sequence whose sum is: \[S_n=\sum_{k=0}^{n-1}0.35^k=\frac{ 1-0.35^n}{ 1-0.35}=\frac{ 1-0.35^n }{ 0.65}\]Then:\[A_n=500·S_n=500·\frac{ 1-0.35^n }{ 0.65 }\]

OpenStudy (anonymous):

After third dose, it should say: \[A_3=A_2⋅(1−0.65)+500=500⋅(1−0.65)^2+500⋅(1−0.65)+500\]

OpenStudy (anonymous):

thank you @CarlosGP

OpenStudy (anonymous):

you are welcome